OpenStudy (anonymous):

Given the functions f(n) = 25 and g(n) = 3(n − 1), combine them to create an arithmetic sequence, an, and solve for the 12th term.

OpenStudy (anonymous):

@ivettef365

OpenStudy (anonymous):

an = 25 − 3(n − 1); a12 = −11 an = 25 − 3(n − 1); a12 = −8 an = 25 + 3(n − 1); a12 = 58 an = 25 + 3(n − 1); a12 = 61

OpenStudy (anonymous):

@ivettef365

OpenStudy (anonymous):

@satellite73

OpenStudy (ivettef365):

If I have to guess I would go for the 3rd option

OpenStudy (anonymous):

thanksss :)

OpenStudy (anonymous):

Given an arithmetic sequence in the table below, create the explicit formula and list any restrictions to the domain.

OpenStudy (ivettef365):

ok, on 4, you have that term 1 = 40, the other terms are greater so n>=40

OpenStudy (ivettef365):

now if you plug term 2, on option 1which is an = 40 + 7(n − 1) let see if it works, term 2 40 + 7(2-1) 40 + 7(1) 40 + 7 = 47 that is correct term 2 is 47

OpenStudy (ivettef365):

did you understand ?

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

So the full answer would then be an = 40 + 7(n − 1) where n ≥ 1 (just took the test!). :) Points earned on this question: 2